trinmjack:

The total number of degrees, D, in an n-sided polygon is given by the formula D = 180 (n minus 2) When Janelle solves for n, she gets n = StartFraction D + 360 Over 180 EndFraction. Sophie solved the same equation for n and got an equivalent equation. Which of the following could be Sophie’s equation?

supie:

It's hard to read this because it says it in words but I believe we would have to do \[d=180(n−2)\] Flip equation \[180n−360=d \] +360 to both sides \[180n−360+360=d+360\]\[180n=d+360\] ÷ both sides by 180 (fraction bar [\(\frac{Blank}{Blank}\)]=division `÷` \(\frac{180n}{180}=\frac{d+360}{180}\) \(n=\frac{1}{180}d+2\) I could be wrong tho...